4 quarts of an alchol solution was mixed with 6 quarts of pure water to make an 8% alchol solution. FInd the percent concentration of the 1st concentraion. you need to show the table.

Answer by stanbon(75887)

4 quarts of an alcohol solution was mixed with 6 quarts of pure water to make an 8% alcohol solution. FInd the percent concentration of the 1st concentration.
------
Equation::
alcohol + alcohol = alcohol
4x + 0*6 = 0.08(x+6)
------
400x + 0 = 8x + 48
392x = 48
x = 12.25% (Percent alcohol in the 1st concentration)
-----------------------
Cheers,
Stan H.
------------